mhchen:

Prove the quotient rule using chain-rule and the derivative of 1/x

myeyeshurt:

You aren't allowed the product rule I assume

myeyeshurt:

@mhchen

myeyeshurt:

Then it's trivial

myeyeshurt:

d/dx u(x)/v(x)=d/dx u(x) * (1/v(x)) = u' *1/v + u*(-1/v^2) and you can do the last algebraic step

myeyeshurt:

And stop thinking in terms of limits that's like trying to do multiplication with repeated addition

myeyeshurt:

Unless you have to use limits @mhchen

myeyeshurt:

In that case expand each derivative into a difference quotient and show the expression holds

myeyeshurt:

Otherwise you can show the product and chain rule hold as well as the derivative of 1/u

myeyeshurt:

Then, redefine the derivative notation as the difference quotient, and just finish what I got you to.

mhchen:

Derivative of 1/x is this: \[\lim_{x \rightarrow c}\frac{\frac{1}{x}-\frac{1}{c}}{x-c} = \lim_{x \rightarrow c}\frac{c-x}{xc(x-c)}\] Now then, derivative of \[(\frac{f(c)}{g(c)})' = f'(c)\frac{1}{g(c)}+f(c)(\frac{1}{g(c)})'\] \[\lim_{x \rightarrow c}\frac{f(x)-f(c)}{(x-c)(g(c)} + f(c)\lim_{x \rightarrow c}\frac{c-g(c)}{g(c)c(g(c)-c)}*\lim_{x \rightarrow c}\frac{g(x)-g(c)}{x-c}\] I used chain rule here \[\lim_{x \rightarrow c} (\frac{(f(x)-f(c))g(c)}{(x-c)g(c)^{2}} - f(c)\frac{g(x)-g(c)}{g(c)c(x-c)})\] \[\frac{\lim_{x \rightarrow c}(\frac{f(x)-f(c)}{x-c})g(c)}{g(c)^{2}}-\frac{f(c)\lim_{x \rightarrow c}(\frac{g(x)-g(c)}{x-c})}{g(c)c}\] My entire work using limits, product rule, and chain rule is here

myeyeshurt:

I can't follow that. Just rewrite each step of my work as a limit and you can show that it holds. Simplify intermediate steps as I did so you don't get limit spaghetti

myeyeshurt:

Also your class Prof is a goofball they might as well make you derive the formula via the epsilon Delta limit definition -.-

justjm:

I never dared to try to understand the epsilon delta limit definition or the 'official definition'

justjm:

Do you need to use limits or can you compose something through if you know what I mean?

myeyeshurt:

The epsilon Delta def is only one way to define it. It's popular in real analysis but not really necessary and impedes understanding when you think about a lot of limits... It's like trying to do counting with the peano axioms

myeyeshurt:

Also @mhchen I'd write out the Sol for you but I'm on mobile and typing is already hard enough